OpenStudy (anonymous):

segment PR has an endpoint at (25, -5) and a midpoint of (18, -1). What is the value of the x-coordinate of the other endpoint?

OpenStudy (anonymous):

do you know the midpoint formula?

OpenStudy (anonymous):

\[\huge\frac{ x_1 + x2 }{ 2 } ~,~ \frac{ y_1+y_2 }{ 2 }\]

OpenStudy (anonymous):

you already know what the midpoint is, so plug the first end point you have and equal it to the midpoint

OpenStudy (anonymous):

I got (21.5,-3) but it was wrong, so I am just trying to figure out the right answer.

OpenStudy (anonymous):

\[\huge\frac{ 25+x }{ 2 } = 18 \]

OpenStudy (shamil98):

You are given the midpoint, and one of the end points.

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

x=11

OpenStudy (mertsj):

From 25 to 18 is 7 units to the left. To get to the other end, go 7 more units to the left. 18-7=11.

OpenStudy (anonymous):

that's correct

OpenStudy (shamil98):

You are asked to find the other endpoint . So plug in your given values into the midpoint formula: midpoint: (18,-1) Endpoint: (25,-5) You need to find the other x value so: \[\huge \frac{ x_1+x_2 }{ 2} = midpoint~coordinate~for~x\] \[\huge \frac{ 25 + x }{ 2 } = 18\] and go from there,

OpenStudy (anonymous):

Thank you!